Find the APR, or stated rate, in each of the following cases (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.):

Stated Rate (APR) Number of Times Compounded Effective Rate (EAR)
% Semiannually 11.0 %
     Monthly 11.9
     Weekly 9.6
     Infinite 13.3
Answer #1

a.EAR=[(1+APR/m)^m]-1
where m=compounding periods

0.11=[(1+APR/2)^2]-1

(1+0.11)=[(1+APR/2)^2]

APR=[(1+0.11)^(1/2)-1]*2

=10.71%(Approx).

b.EAR=[(1+APR/m)^m]-1
where m=compounding periods

0.119=[(1+APR/12)^12]-1

(1+0.119)=[(1+APR/12)^12]

APR=[(1+0.119)^(1/12)-1]*12

=11.3%(Approx).

c.EAR=[(1+APR/m)^m]-1
where m=compounding periods

0.096=[(1+APR/52)^52]-1

(1+0.096)=[(1+APR/52)^52]

APR=[(1+0.096)^(1/52)-1]*52

=9.17%(Approx).

d.EAR=(e)^APR-1

where e=2.71828

0.133=(2.71828)^APR-1

(1+0.133)=2.71828^APR

Taking log on both sides;

log 1.133=APR*log 2.71828

APR=log 1.133/log 2.71828

=12.49%(Approx).